# Tactile SoftHand-A: 3D-Printed, Tactile, Highly-underactuated, Anthropomorphic Robot Hand with an Antagonistic Tendon Mechanism

### v1

@@ 8 Results
8.1 Results A - Finger Gestures Comparison
We use the D-type finger as an example to illustrate the active control of the DIP joint through coordinated control of the driving and antagonistic motors. Initially, in the finger extension state, the driving motor is position-controlled to rotate a specific angle, pulling the driving tendon and generating a driving force on the finger. Concurrently, the active antagonism is position-controlled to slightly reverse, thus eliminating the slack in the tendon. As a result, substantial opposing forces are generated at the PIP and MCP joints. Due to the driving tendon passing through the entire finger and resistance at the MCP and PIP joints, active control of the DIP joint is achieved. Figure 8 D1(a) displays the gestures of the D-type finger in actual experiments. The green line in Figure 8 D1(b) represents the changing angle of the DIP, while Figure 8 Input 1 details the input changes to the driving and antagonistic motors in this procedure. Figure 8D1(a-b) shows the D-type finger achieving active control of the DIP in a semi-flexion state, following a similar control process. Figure 8D3(a-b) shows the D-type finger in full flexion. The control of the P-type and A-type fingers are similar to that of the D-type, with detailed results in Figure 8.
8.2 Results B - Controllable Grasping Gestures
As illustrated in Figure 9A(a), under the open hand gesture of the Tactile SoftHand-A, control over all fingers’ DIP joints is achieved by increasing the differential between the driving motor and the antagonistic motor. Notably, in the initial phase, increasing the encoder of the antagonistic motor (to reverse its direction at phase 3) tightens the antagonistic tendon, thereby preventing the fingers’ PIP and MCP joints from bending actively. The actual experiment’s representations, including the front and side views of the Tactile SoftHand-A, are displayed in Figure 9B(d-e). Here, the bending of all fingers’ DIP joints is distinctly observable as the motor differential increases, reaching its maximum at phase 3.
In Figure 9A(b), we first reduce the encoders of both the driving motor and the antagonistic motor simultaneously to coordinate their rotation (phase 1-2), and then increase the differential between them (phase 3-4). This enables active control of the DIP joints while the hand is in a semi-closed state. The corresponding actual presentations are depicted in Figure 9B(f-g). We observe that in phases 1-2, the Tactile SoftHand-A begins to close. Subsequently, in phase 3, active control over the rotation of the fingers’ DIP joints commences, reaching its maximum in phase 4, followed by a return to an open position in phase 5.
By contrast, in Figure 9A(c), coordinating the rotation of the driving motor and the antagonistic motor without activating the active antagonism allows for the hand to close and open normally.
8.3 Results C - Grasping Adaptability
@@
From Figures 10A and B, we observe the interaction between the Tactile SoftHand-A and the grasped objects. Specifically, Figure 10A(a) illustrates how the thumb, positioned opposite the other four fingers, forms an arc to effectively encircle and stabilize objects, enhancing grip support. This arrangement and differential mechanism allow the remaining fingers to flex to various extents, adapting to the object’s shape and demonstrating the Tactile SoftHand-A’s adaptability to different object contours.
In the case of grasping single-section objects include the hexagonal prism, triangular prism, and cylinder, the fingers, upon making contact, permit the little finger to retract towards the palm, as depicted in Figure 10B(a-c). This flexibility is attributed to the novel differential mechanism, which is not limited to the little finger alone.
@@
Furthermore, as shown in Figure 10A and B(d-e), when engaging with multi-section objects, each finger joint adjusts its bending to conform to the object’s shape, showcasing the Tactile SoftHand-A’s adaptability in grasping varied geometries. For example, after the little finger, thumb, and index finger of the Tactile SoftHand-A make contact with the outer flange of the tendon spool, the ring and middle fingers can still bend until they touch the central axis of the tendon spool.
@@ 8.4 Results D - Tactile Feedback Control
Figure 13 depicts various time points during the gesture synchronization phase (1-15 seconds), where 1 to 3 represent specific moments of synchronization stage. At this stage, Tactile SoftHand-A synchronizes with the gestures of the human hand. For instance, at time point 1, Figure 13(A) illustrates the human hand in an open position, while Figure 13(B) shows Tactile SoftHand-A in a similar fully open state. By time point 3, Figure 13(A) depicts the human hand partially closed, and the Tactile SoftHand-A mirroring this gesture in Figure 13(B). Figure 13(C)(2) demonstrates that the input curves of the two motors follow a similar trend to the finger angle curve of the human hand, showcasing effective synchronization of the human and robot hand gestures. Notably, the delay between these curves is less than one second, attributable to the anti-tendon mechanism in Tactile SoftHand-A, which facilitates rapid response.
@@
From approximately 26 to 33 seconds, the gesture synchronization phase occurs, during which Tactile SoftHand-A is controlled by the human hand to re-establish contact. From 33 to 43 seconds, the slip phase takes place. During this phase, Figure 13(D)(7-8) illustrate the changes in the contact area and the slip of its center point, clearly showing the contact point moving first to the right and then slipping to the left. The green curve in Figure 13(C)(3), which represents slip detection, also remains in the slip state during this period, indicating that slippage has occurred. At same time, the contact is also indicated (blue curve). Following the onset of slippage, the motor inputs change as shown in moments 7 and 8 of Figure 13(C)(2), adjusting the grip gesture of Tactile SoftHand-A by increasing the input differential between the driving motor and the antagonistic motor. As demonstrated in Figure 13(B)(7-8), the DIP joints of Tactile SoftHand-A’s fingers actively rotate inward, adjusting their grip to enhance the gripping force and prevent the object from slipping.
@@ 9 Discussion
The experimental results presented highlight the intricate and precise control mechanisms implemented in the Tactile SoftHand-A, in particular with the D-type finger permitting control of the distal interparietal joint. The key findings from these results are discussed below.
The D-type finger experiments illustrate the effectiveness of coordinated control between the driving and antagonistic motors. By managing the position and tension of the tendons, suitable forces are generated at the PIP and MCP joints that in turn allow for precise control of the DIP joint. This level of control is crucial for tasks requiring fine motor skills, such as changing the hand gesture to prevent slippage while not appreciably increasing the grasping force. The experimental data, depicted in Figure 8D1(a-b), D2(a-b) and D3(a-b), demonstrate the ability of the D-type finger to maintain various degrees of flexion, showing both semi-flexion and full flexion states. The results confirm that the coordinated tendon mechanism effectively controls the finger joints, enhancing the overall dexterity and functionality of the tendon-driven fingers.
The results for the Tactile SoftHand-A in Figure 9 underscore the significance of coordinated motor control in achieving complex hand gestures. By adjusting the encoder values of the driving and antagonistic motors, the DIP joints of all fingers can be controlled, as shown in Figure 9. This method prevents the PIP and MCP joints from active bending, allowing for isolated control of the DIP joints, which is essential for tasks requiring a stable grip while maintaining finger dexterity. The transition from an open hand to a semi-closed and then fully closed hand demonstrates the adaptability of the control system in managing different phases of hand movement.
@@
The adaptability of the Tactile SoftHand-A is further evidenced in grasping experiments with various objects (Figure 10). The thumb’s positioning opposite the other fingers allows for effective encircling and stabilization of objects, enhancing the grip’s support. The differential mechanism enables each finger to adjust its flexion based on the object’s shape, demonstrating a high level of adaptability. This characteristic is particularly important for robot hands that must interact with a variety of objects in daily life. The ability of the fingers to retract and bend independently ensures that the hand can conform to different geometries, improving the practicality and usability of the robot hand.
@@
The tactile feedback mechanism in the Tactile SoftHand-A (Figure 7) highlights the system’s ability to synchronize gestures with a human hand and respond to external stimuli. The synchronization phase demonstrates effective mirroring of human hand movements, with a minimal delay attributed to the antagonistic tendon mechanism. The contact and slip detection phases evidence the hand’s ability to maintain a stable grasp and adjust to prevent slippage. This closed-loop tactile feedback control is crucial for tasks that require a secure grip, such as holding delicate objects or performing precision tasks. The ability to detect and respond to slippage ensures that the robot hand can adapt to dynamic conditions, enhancing its functionality and reliability.

### v2

@@ 8 Results
8.1 Experiment A Results
8.1.1 Results A1 - Comparison of Finger Gestures:
First, we use the D-type (distal-type) finger to illustrate the decoupled control of the DIP joint through coordinated control of the agonist and antagonist motors. Initially, in the finger extension state, the agonist motor is position-controlled to rotate a specific angle, pulling the agonist tendon and generating a closing force on the finger. The antagonist motor can be position controlled to modify the opposing tension of the antagonist tendon, giving different gestures of the finger (motor inputs 1, 2 and 3 in Figure 8), and also to extend the finger to actively return it to an open posture.
The first motor input flexes the finger under the agonist motor (input 1 in Figure 8) while allowing a large differential with the antagonist motor. The D-type and P-type fingers differ in that the D-type can be controlled to flex just the DIP joint due to opposing forces at the PIP and MCP joints, unlike the P-type where the MCP and PIP joints are also flexed. This process is reversed to extend the finger.
The second motor input has two initial stages, starting with a small flexion of the finger in which the active antagonism is reversed to lower the antagonist tension, then a larger drive flexes the entire finger (input 2 in Figure 8). The D-type and P-type fingers behave similarly in flexing the entire finger, except the P-type finger has a larger flexion. Again, this process is reversed to extend the finger.
The third motor input is similar to the first input to flex the finger, but has a larger de-tensioning of the antagonist tendons (input 3 in Figure 8). All three D-type, P-type and A-type fingers behave similarly in flexing all joints.
Figure 11: Grasping adaptivity of the tactile SoftHand-A with D-type fingers to various objects with the same motor inputs. (A) Front and side views of the hand at maximal closure. (B) Input from the agonist and antagonist motors. (C) The number of contacting fingertips at maximum closure (validated from the tactile outputs).
Figure 12: Comparison of effective grasping with with Tactile SoftHand-A (shown in blue) and the BPI SoftHand with a passive reset (shown in red) incorporating passive elements. Results for the BPI SoftHand are from Li et al. (2022).
8.1.2 Results A2 - Dynamic Performance:
Multiple performance measures for a single D-type finger of the SoftHand-A are summarized in Table 1 (with experiments shown in Appendix Figure 17). These tests assessed properties such as response speed, repeatable positioning accuracy, bearing capacity, and manipulation capability. In all cases, the results are compared with a single finger of the BPI SoftHand (Li et al. (2022)) to quantify the performance improvements from active antagonism.
Overall, the D-type finger of the SoftHand-A is able to rapidly flex and extend, taking less than 0.5 sec between fully open and closed positions (corresponding to joint motions of 2-3 rad/sec). In contrast, the finger of the BPI SoftHand is slower to flex to full closure (by 0.17 sec), but can more rapidly extend to opening (by 0.14 sec) due to the passive elastic reset that is activated immediately when tension is removed from the agonist motor.
Likewise, the repeat positioning is more accurate for the D-type finger of the SoftHand-A (within $\pm 0.3$ mm) than for a finger of the BPI SoftHand (within $\pm 0.5$ mm). This is attributed to the more precise control from having both agonist and antagonist motors.
The single finger bearing capacity is much higher for the SoftHand-A finger (8 N) compared tothe BPI SoftHand finger (0.9 N). This is expected because the SoftHand-A is able to bear load by tensioning the antagonist tendon, but the BPI SoftHand only has a passive elastic mechanism to maintain finger extension.
The single finger manipulation experiments compared simple actions such as pushing, pulling and sliding from above (Appendix Figure 17D,E). For a baseline, pushing with the entire finger was considered, with a maximum load of 7 N for the SoftHand-A that was 5 N larger than that for the BPI SoftHand. Finally, only the SoftHand-A could demonstrate fingertip object manipulation, as the position of its DIP joint is controllable (see previous section on grasp gestures). The experimental results showed that by utilizing the two-tendon system, the DIP joint could be actively controlled while simultaneously reactivating the antagonist tendon to rotate the MCP joint, providing sufficient workspace for the fingertip to manipulate the object by pushing, pulling and sliding.
8.2 Experiment B Results
8.2.1 Results B1 - Controllable Grasping Gestures:
The SoftHand-A uses the D-type joint so that control over the DIP joints of all fingers can be achieved by increasing the differential between the agonist motor and the antagonist motor. Initially, activating the antagonist motor tightens the antagonist tendon, thereby preventing the fingers’ PIP and MCP joints from bending actively; then its direction can be reversed in phase 3 to change the DIP joint angles. The experiments show the front and side views of the SoftHand-A under three motor inputs analogous to those used to assess single finger gestures (Figures 9).
The first motor input flexes the fingertips using the agonist motor with a large differential with the antagonist motor (input 1 in Figure 9). This motor activation flexes just the DIP joints due to opposing forces at the PIP and MCP joints. The reverse process extends the fingertips by rotating just the DIP joints.
The second motor input has an initial stage that starts with a small flexion of the finger in which the active antagonism is reversed to lower the antagonist tension, followed by a larger drive that closes the entire hand (input 2 in Figure 9). The initial stage then enables active control of the DIP joints while the hand is in a semi-closed state (Figure 9, middle panels). The hand is opened by activating the agonist and antagonist motors to return to their initial positions.
The third motor input is similar to the first input to flex the finger, but drives the antagonist motor to decrease the differential between the two motors, de-tensioning the antagonist tendons (input 3 in Figure 9). This motion also results in a grasp gesture that closes the entire hand, but without the active control of just the DIP joints. The hand is opened by returning the motors to their initial positions.
Figure 13: Reaction to contact disturbances with the Tactile SoftHand-A. (A) Interactions between Tactile SoftHand-A and indenter, showing 10 stages where the hand closes onto the indenter, the the fingertip is pushed, pulled (by withdrawing the indenter) and pushed back again. The red arrows show the direction that the indenter is moving. (B) Calibration of tactile deformation (1-SSIM) with normal force measured from a strain gauge, where SSIM measures the similarity of the tactile image with a reference undeformed image. (C) Top, change in inferred force over time for the tactile fingertip of the index finger; bottom: change in encoder values over time for the agonist and antagonist motors.
Figure 14: Human hand gesture and tactile control of the Tactile SoftHand-A. (A) human hand gestures used as inputs to control the hand. (B) Side view of Tactile SoftHand-A in response to the gesture control. (C-1) Average degree of human finger closure, (C-2) corresponding motor inputs to the hand, (C-3) and change of finger contact and slip. (D) Tactile contact map at various stages throughout the motion as the hand contacts and object. The results in this figure relate to Experiment D.
Figure 15: Comparison of grasping under an external disturbance with and without tactile feedback (left and right panels, respectively). Phase 1: a teleoperator closes the hand to stably hold the object. Phase 2: an external disturbance is applied to the held object. Phase 3: with no tactile feedback the hand is unable to respond to this disturbance, otherwise with tactile feedback the hand can tighten its grip to prevent the object being dropped.
8.2.2 Results B2 - Hand Performance:
The hand performance evaluation is summarized in Table 2 and experiments depicted in the Appendix (Figure 17). Overall, the hand demonstrated rapid movements of $\sim 0.5$ sec between fully open and closed positions with a maximum normal wrench force of 20 N and 9 N in the other directions, and maximum wrench torques of $0.3$ Nm to $0.6$ Nm.
A more detailed depiction of the wrench forces and torques over time is shown in the Appendix (Figure 17(B)). The forces $(F_{x},F_{y},F_{z})$ are applied within the coordinate frame of the hand. We note that the normal force $F_{z}$ peaks at around 20 N. Similarly, the maximal torque is in the axis $R_{z}$ and reaches a peak value of 0.6 Nm. These plots illustrate the robotic hand’s capacity to apply significant forces and torques, indicating its capability for physical interaction during various tasks.
The SoftHand-A is able to effectively form power, fingertip and pinch grips on various objects (Figure 10). When a power grasp is employed, the fingers conform to the surface of the object, increasing the contact area and allowing for a secure grasp (Figure 10A), which is particularly suitable for grasping heavier objects. For delicate objects with easily damaged surfaces, such as fruits, the SoftHand-A can switch to the Fingertip Grasp mode (Figure 10B), like the human hand. In this mode, only the soft fingertips make contact with the object’s surface, effectively avoiding the potential for surface damage that may occur in power grasp mode due to the uncontrolled rotation or pressure across the hand. Finally, the SoftHand-A is also able to grasp a range of small or thin objects with pinch grasps between the thump and forefinger (Figure 10c). This flexibility in grasping modes ensures that the Tactile SoftHand-A can handle a wide variety of objects while minimizing the risk of damage to sensitive items.
8.3 Results C - Grasping Adaptivity
@@
The interactions between the Tactile SoftHand-A and various grasped objects are shown in Figures 12A. Specifically, the thumb, positioned opposite the other four fingers, forms an arc to encircle and stabilize objects. This arrangement and differential mechanism allow the remaining fingers to flex to various extents, adapting to the object’s shape, which demonstrates the Tactile SoftHand-A’s adaptivity to different object geometries.
For many of the objects, the little finger (pinkie) to retracts towards the palm, giving additional enclosure, as is most evident in Figures 12a5-a8. This adaptivity is attributed to the novel differential mechanism, which is not limited to the little finger alone, and is also evident in the variation in the number of contracted fingertips (Figure 12C).
@@
For the irregular objects, each finger joint adjusts its bending to conform to the object’s shape (see Figures 12a1-a6), showcasing the SoftHand-A’s adaptability in grasping varied geometries. For example, after the little finger, thumb, and index finger of the Tactile SoftHand-A make contact with the outer flange of the tendon spool, the ring and middle fingers can still bend until they touch the central axis.
In Figure 12(D), we compared the grasping ability of the Tactile SoftHand-A with that of the BPI SoftHand incorporating passive elements (Li et al. (2022)). For most objects, the Tactile SoftHand-A demonstrated superior or equivalent performance, successfully grasping many items in five out of five attempts. Objects such as cuboids, cylinders, cable spools, and several soft toys (e.g., a strawberry toy, popcorn toy, and orange toy) consistently exhibited higher grasp success rates, with the Tactile SoftHand-A achieving the maximum possible success over five grasps. For some smaller or irregularly shaped objects, such as a ball, pen and coins, or heavier objects like detergent, the success rate dropped for both hands, indicating the difficulty these hands face with these items. However, the Tactile SoftHand-A generally maintained higher consistency and adaptivity across a wider range of tested objects, especially with everyday tools and toys. This comparison highlights the enhanced robustness and effectiveness of the Tactile SoftHand-A in handling diverse shapes and sizes, particularly in scenarios requiring more precise or adaptive grasping strategies.
@@ 8.4 Results D - Tactile Feedback Control
8.4.1 Results D1 - Reaction to Contact Disturbances:
Various responses of the Tactile SoftHand-A during changes in contact are shown in Figure 13(A). Throughout the experiment, the contact (here a tip of a pen) underwent a series of actions implemented manually, such as being pushed forward or retracted (phases 0-9). In the initial phases (phases 1-2), the tactile fingertips were not yet in contact with the pen, and the Tactile SoftHand-A closed its fingers until contact was made in phase 3. Subsequently, the pen was manually moved upward, resulting in increased pressure on the tactile fingertips and the Tactile SoftHand-A responded by extending its fingers to maintain a target deformation and preserve a stable contact. In phases 6-7, as the pen continued to move upward, the fingers extended further in response to this motion. Then in phase 8, when the pen began to retract, the fingers instead flexed to maintain the deformation.
As described in the experimental sections, the control law seeks to maintain a fraction of deformation value at a target of 0.05, quantified by the structural similarity index measure (SSIM) of the current tactile image from an undeformed tactile image. A calibration of 1-SSIM against normal force (Figure 13(B)) revealed a near-linear relation, such that the target corresponds to a normal force of approximately 2 N. Consequently, the control law seeks to adjust the motor encoder values to maintain a normal force of this value, which is evident in the plots of these quantities over time (Figure 13(C)).
These experimental results demonstrate that the Tactile SoftHand-A can use its tactile feedback to effectively maintain a stable contact during dynamic tasks.
8.4.2 Results D2 - Reactive Teleoperation:
This experiment is structured into three sequential stages: synchronization, contact, and slip. In the synchronization stage, the Tactile SoftHand-A aligns with the gestures of the human hand. During the contact and slip stages, the system engages in closed-loop tactile feedback control. The results are shown in Figure 14.
Various time points during the gesture synchronization phase (1-15 seconds), are shown in Figure 14 from time points 1-3. At this stage, Tactile SoftHand-A synchronizes with the gestures of the human hand. For instance, at time point 1, Figure 14(A) illustrates the human hand in an open position, while Figure 14(B) shows the Tactile SoftHand-A in a similar fully open state. By time point 3, Figure 14(A) depicts the human hand partially closed, and the Tactile SoftHand-A mirroring this gesture in Figure 14(B). Figure 14(C)(2) demonstrates that the input curves of the two motors follow a similar trend to the finger angle curve of the human hand, showcasing effective synchronization of the human and robot hand gestures. Notably, the delay between these curves is less than one second, attributable to the anti-tendon mechanism in Tactile SoftHand-A, which facilitates rapid response.
@@ 8.4.2 Results D2 - Reactive Teleoperation:
From approximately 26 to 33 seconds, the gesture synchronization phase occurs, during which Tactile SoftHand-A is controlled by the human hand to re-establish contact. From 33 to 43 seconds, the slip phase takes place. During this phase, Figure 14(D)(7-8) illustrates the changes in the contact area and the slip of its center point, clearly showing the contact point moving first to the right and then slipping to the left. The green curve in Figure 14(C)(3), which represents slip detection, also remains in the slip state during this period, indicating that slippage has occurred. At the same time, the contact is also indicated (blue curve). Following the onset of slippage, the motor inputs change as shown in moments 7 and 8 of Figure 14(C)(2), adjusting the grip gesture of Tactile SoftHand-A by increasing the input differential between the agonist motor and the antagonist motor. As demonstrated in Figure 14(B)(7-8), the DIP joints of Tactile SoftHand-A’s fingers actively rotate inward, adjusting their grip to enhance the gripping force and prevent the object from slipping.
8.4.3 Results D3 - Reaction to Object Disturbances:
Our final experiment extends the teleoperated control to grasping performance when external disturbances are applied to the held object, comparing the influence with and without tactile feedback (see Figure 15).
In the non-tactile feedback experiments, the SoftHand-A first successfully grasped a soft toy (shaped like a strawberry) using a teleoperation procedure (denoted Phase 1), with inferred fingertip contact forces stabilizing around 3.5 N. When external disturbances were applied (Phase 2), the inferred contact force rapidly increased to $\sim$8 N, followed by an abrupt decline to 0 N (denoted Phase 3) due to object slippage (Figure 15A). The entire experiment was repeated with a stiffer object (a cup), which equilibrated to a higher contact force of 12 N, before likewise failing to securely hold the object under an external disturbance (Figure 15C).
In contrast, the experiments with tactile feedback exhibited successful reactions to external disturbances, as the hand could react autonomously to stabilize the object. During Phases 1 and 2, the inferred contact forces remained consistent with the non-tactile experiments. However, phase 3 was significantly different: upon detecting a change in contact force induced by disturbances, the control system triggered a compensation mechanism that adjusted just the DIP joint to bring the fingertips into greater contact with the object. While maintaining object retention and counteracting dynamic external disturbances, the Tactile SoftHand-A achieved a reduction of contact forces, ultimately establishing an equilibrium between the disturbance components of the force and the other contact forces to achieve a stable grasp. This tactile feedback successfully prevented secondary object slippage, in contrast to the object loss observed without tactile feedback.
These comparative experiments demonstrate the adaptive control capabilities of using tactile feedback coupled with the antagonistic tendon mechanism of the tactile SoftHand-A.
@@ 9 Discussion
In this paper, we have presented and tested our design of the Tactile SoftHand-A, a 3D-printed highly-underactuated, anthropomorphic robot hand with an antagonist tendon mechanism based on the BPI-SoftHand (Li et al. (2022)), a 3D-printed derivative of the Pisa/IIT SoftHand (Catalano et al. (2014)). The SoftHand-A has improved dexterity over the original purely agonist mechanisms with passive reset through elastic elements. The Tactile SoftHand-A also features vision-based tactile sensors that are completely 3D-printed within the hand structure, using the TacTip design of marker-based tactile sensors (Ward-Cherrier et al. (2018); Lepora (2021)). The experimental results assessed the novel control capabilities possible with the Tactile SoftHand-A. The key findings from these results are discussed below.
First, we tested the performance of a single finger, considering three versions differing in the termination of the antagonist tendon: the A-type finger terminating at the tip of the distal phalange, the D-type finger on the medial phalange, and the P-type on the proximal phalange (Figure 8). Of these, the A-type finger is the most reactive; however, the D-type finger permits control of the distal interphalangeal (DIP) joint. By managing the position and tension of the tendons, suitable forces are generated at the PIP and MCP joints to allow precise control of the DIP joint. This level of control enables tasks requiring fine motor skills, such as changing the hand gesture to prevent slippage while not appreciably increasing the grasping force. Because of this versatility, this design is adopted in the SoftHand-A.
Quantitative comparisons between a single finger of the SoftHand-A (D-type) and the BIP SoftHand showed several benefits of this design. Overall, the antagonistic mechanism results in a moderate improvement in reaction speed and accuracy, and a large increase in weight bearing capacity (Table 1). Overall, the most significant improvements are in manipulation capabilities such as pushing, pulling, and sliding, which the precursor BPI SoftHand is unable to perform.
Next, we tested the whole-hand performance of the SoftHand-A, considering the version with the D-type finger due to its improved manipulation capabilities. By adjusting the agonist and antagonist motors, the DIP joints of all fingers can be controlled, as shown in Figure 9. This method prevents the PIP and MCP joints from active bending, allowing for isolated control of the DIP joints, which is needed for tasks requiring a stable grip while maintaining finger dexterity. As a consequence, the Tactile SoftHand-A is capable of performing various grasping modes, including power grasp, fingertip grasp, and pinch grasp that all benefit from the hand’s adaptivity onto various object shapes (Figure 10). The fingertip grasp is a new grasp enabled by the D-type finger that is suitable for handling delicate objects, where reduced contact between the hand and the object protects a fragile surface while maintaining a stable grip. By switching between these different modes, the SoftHand-A can flexibly and adaptively adjust its grasping strategy based on the characteristics of the object.
A quantitative comparison of the Tactile SoftHand-A and the BPI SoftHand shows significantly improved reaction times and wrench forces/torques due to the antagonistic tendon mechanism (Table 2). Improvements are around a factor of two, with the closure time differing most (reducing from 1.4 sec to 0.5 sec). These originate from not having to work against a restoring force that the BPI SoftHand uses to passively reset to an open position.
@@
The adaptivity of the Tactile SoftHand-A is further evidenced in grasping experiments with various objects (Figure 12). The thumb’s positioning opposite the other fingers allows for effective encircling and stabilization of objects, enhancing the grip’s support. The differential mechanism enables each finger to adjust its flexion based on the object’s shape, which is particularly important for robot hands that must interact with a variety of objects in daily life. Likewise, a comparison with the BPI SoftHand demonstrated a significantly improved adaptive capability to achieve secure grasps (Figure 12).
Then three experiments test the sensing capabilities of the Tactile SoftHand-A and the reactive control that the tactile sensing gives. First, we test the reactivity to contact disturbances through real-time feedback from the tactile fingertips (Figure 13). A feedback controller based on stabilising the tactile deformation (estimated from the SSIM and equivalent to contact force) dynamically adjusts the finger posture via activating the agonist and antagonist motors to ensure stable contact. During dynamic contact interactions, continuous variations in contact force and direction between the indenter and tactile fingertip induce changes in the tactile images, from which the Tactile SoftHand-A adaptively regulates its fingertip positions to preserve a preset contact threshold. This demonstrates the advantage of tactile sensing in counteracting contact disturbances.
@@
The second experiment demonstrates the tactile feedback mechanism in the Tactile SoftHand-A (Figure 7) to synchronize gestures with a human hand and respond to external stimuli. The synchronization phase demonstrates effective mirroring of human hand movements, with a minimal delay attributed to the antagonistic tendon mechanism, which reduces reliance on passive elements within the finger joints, thereby enhancing responsiveness. Furthermore, the contact and slip detection phases demonstrate the Tactile SoftHand-A’s ability to maintain a stable grasp while dynamically adjusting its gesture to prevent slippage (Figure 14).
Finally, the third experiment comparatively analyzes the performance of tactile slip detection and grasp gesture compensation mechanisms using the teleoperation of the second experiment (Figure. 15). Under disturbance-free conditions, both tactile-enabled and tactile-deprived grasping modalities achieve stable object retention. However, when dynamic disturbances were applied to the manipulated object, the non-tactile system fails to detect grasp failure (slip events), ultimately resulting in object loss. In contrast, the tactile-enabled Tactile SoftHand-A successfully demonstrates detecting slip occurrences and triggers compensatory grasp adjustments, thereby maintaining a stable grasp. This closed-loop tactile feedback control is crucial for tasks that require a secure grip, such as holding delicate objects or performing precision tasks. By effectively detecting and responding to slippage, the robotic hand can adjust to dynamic environmental conditions, significantly improving its functionality, reliability, and overall performance in real-world applications.
Overall, these three experiments using tactile feedback confirm the critical role of tactile perception in enhancing disturbance resistance and improving grasping reliability within unstructured environments, and how this complements the adaptivity and controllability of the SoftHand-A due to its antagonistic tendons and synergistic mechanism.